# protoplanetary disks — planets and satellites: composition

### v1

@@ 3.2.1 Protoplanet formation time
Figure 7: Same as the right panel of Figure 6 but from the simulations with protoplanet formation ages of 0.1 Myr (Case 1, left panel) and 1.0 Myr (Case 2, right panel).
@@ 3.2.2 Core orbit
Figure 8: Same as the right panel of Figure 6 from the simulations with core orbits of 5.2 au (Case 3, left panel) and 10.0 au (Case 4, right panel).
@@ 3.2.3 Dissociation temperature
Figure 9: Same as the right panel of Figure 6 but from the model where the salt is assumed to be NH4CN, with $T_{\rm NH_{4}CN,vap}=150$ K and $f_{\rm NH_{3},NH_{4}CN}=0.4$ (Case 5).
@@ 3.2.4 Fragmentation velocity
Figure 10: Same as Figure 4 and the right panel of Figure 5 but from the model with $v_{\rm frag}=1.0~{}\rm m~{}s^{-1}$ (Case 6).
Figure 11: Same as the right panel of Figure 6 but from the model with $v_{\rm frag}=1.0~{}\rm m~{}s^{-1}$ (Case 6).
@@ 3.2.5 Turbulence strength
Figure 12: Same as Figure 4 and the right panel of Figure 5 but $\alpha=10^{-4}$ (Case 7).
Figure 13: Same as the right panel of Figure 6 but $\alpha=10^{-4}$ (Case 7).
@@ 3.2.6 Effect of recondensation
Figure 14: Same as Figure 4 and the right panel of Figure 5 but this model included the effect of recondensation (Case 8).
Figure 15: Same as the right panel of Figure 6 but this model included the effect of recondensation (Case 8).
@@ 5 Summary and Conclusions
We would like to thank Kenji Furuya, Kazumasa Ohno, and Shota Notsu for useful discussions. We also thank Kazuaki Homma for advice on our numerical simulations. We are grateful to the anonymous referee for helpful comments.
6 Funding
This work was supported by JSPS KAKENHI Grant Numbers JP20H00205, JP23H01227, JP23H00143, and JP23KJ0907.
@@ Appendix B Elements entrapment by amorphous ice
The difference in entrapment efficiency can be larger when the target is a multi-component gas mixture, due to intermolecular competition for adsorption onto the matrix, called competitive entrapment. Bar-Nun et al. (2007) examined the entrapment efficiencies for a gas mixture of CO, N2, and Ar, and reported that the entrapment efficiency for N2 is more than an order of magnitude lower than those for CO and Ar. The low entrapment efficiency for N2 is also supported by the fact that comets’ N2/CO ratio is lower than that of the protosolar nebula value (Anderson et al., 2023).

### v2

@@ 3.2.1 Protoplanet formation time
Figure 7: Same as the right panel of Figure 6 but from the simulations with protoplanet formation ages of 0.1 Myr (Case 1, left panel) and 1.0 Myr (Case 2, right panel). Alt text: Two line graphs showing the evolution of atmospheric nitrogen abundance for different core insertion times.
@@ 3.2.2 Core orbit
Figure 8: Same as the right panel of Figure 6 from the simulations with core orbits of 5.2 au (Case 3, left panel) and 10.0 au (Case 4, right panel). Alt text: Two line graphs showing the evolution of atmospheric nitrogen abundance for different core orbits.
@@ 3.2.3 Dissociation temperature
Figure 9: Same as the right panel of Figure 6 but from the model where the salt is assumed to be NH4CN, with $T_{\rm NH_{4}CN,vap}=150$ K and $f_{\rm NH_{3},NH_{4}CN}=0.4$ (Case 5). Alt text: Line graph showing the evolution of atmospheric nitrogen abundance when the anion of ammonium salt differs.
@@ 3.2.4 Fragmentation velocity
Figure 10: Same as Figure 4 and the right panel of Figure 5 but from the model with $v_{\rm frag}=1.0~{}\rm m~{}s^{-1}$ (Case 6). Alt text: Two line graphs showing the distribution of NH3 molecules before and after core insertion for the case of low dust stickiness.
Figure 11: Same as the right panel of Figure 6 but from the model with $v_{\rm frag}=1.0~{}\rm m~{}s^{-1}$ (Case 6). Alt text: Line graph showing the evolution of atmospheric nitrogen abundance for the case of low dust stickiness.
@@ 3.2.5 Turbulence strength
Figure 12: Same as Figure 4 and the right panel of Figure 5 but $\alpha=10^{-4}$ (Case 7). Alt text: Two line graphs showing the distribution of NH3 molecules before and after core insertion for the case of weak turbulence.
Figure 13: Same as the right panel of Figure 6 but $\alpha=10^{-4}$ (Case 7). Alt text: Line graph showing the evolution of atmospheric nitrogen abundance for the case of weak turbulence.
@@ 3.2.6 Effect of recondensation
Figure 14: Same as Figure 4 and the right panel of Figure 5 but this model included the effect of recondensation (Case 8). Alt text: Two line graphs showing the distribution of NH3 molecules before and after core insertion for the model considering recondensation.
Figure 15: Same as the right panel of Figure 6 but this model included the effect of recondensation (Case 8). Alt text: Line graph showing the evolution of atmospheric nitrogen abundance for the model considering recondensation.
@@ 4.6 Assumptions and limitations in the model
Core dissolution degree: Core dissolution plays a crucial role in explaining Jupiter’s nitrogen enrichment, except in cases where a low dissociation temperature (Figure 9), weak turbulence (Figure 13), and the effect of recondensation are considered (Figure 15), allowing most of Jupiter’s nitrogen enrichment to be explained solely by gas accretion. In our model, we have assumed that 10 $M_{\rm\oplus}$ (0.03 $M_{\rm Jup}$) of heavy elements dissolve from the core into the atmosphere. If the mass fraction of salt in the dust is 20 wt%, this process would result in a nitrogen abundance in Jupiter’s atmosphere of $\sim$ 1.4 times the solar. Assuming the mass fraction and type of salt remain constant, the degree of nitrogen enrichment due to core dissolution is proportional to the dissolved core mass. Juno observations constrain the total heavy-element mass in Jupiter to 6–27 $M_{\rm\oplus}$, implying that core dissolution could increase the nitrogen abundance in the atmosphere by up to a factor of 3.8.
@@ 5 Summary and Conclusions
We would like to thank Kenji Furuya, Kazumasa Ohno, and Shota Notsu for useful discussions. We also thank Kazuaki Homma for advice on our numerical simulations. We are grateful to the anonymous referee for helpful comments. This work was supported by JSPS KAKENHI Grant Numbers JP20H00205, JP23H01227, JP23H00143, and JP23KJ0907.
@@ Appendix B Elements entrapment by amorphous ice
The difference in entrapment efficiency can be larger when the target is a multi-component gas mixture, due to intermolecular competition for adsorption onto the matrix, called competitive entrapment. Bar-Nun et al. (2007) examined the entrapment efficiencies for a gas mixture of CO, N2, and Ar, and reported that the entrapment efficiency for N2 is more than an order of magnitude lower than those for CO and Ar. Zhou et al. (2024) re-evaluated the effect of competitive entrapment on mixtures of multicomponent gases, including CO, CH4, N2, and Ar. Their experimental results confirmed that the entrapment efficiency of N2 is lower by a factor of two to an order of magnitude compared to other molecules when the deposition temperature exceeds 40 K. However, they did not reproduce the decrease in N2 entrapment efficiency due to competitive entrapment at lower temperatures, as reported by Bar-Nun et al. (2007). They suggested that factors such as the H2O/target ratio and the deposition rate of amorphous ice may influence the competitive entrapment process. The low entrapment efficiency for N2 is also supported by the fact that comets’ N2/CO ratio is lower than that of the protosolar nebula value (Anderson et al., 2023). Zhou et al. (2024) proposed that this low N2/CO may originate from amorphous ice that formed at 40–50 K in the protosolar nebula.